# Enhancing the Sensitivity to Seesaw Predictions in Gauged $B-L$ Scenarios

### v2

@@ IV Main results
Figure 7: Background-free sensitivity contours at 90% C.L. for SBND, MicroBooNE, ICARUS-BNB, and DarkQuest as a function of $m_{N}$ and $\left|U_{\mu}\right|^{2}$ (left column) and $\left|U_{\tau}\right|^{2}$ (right column). Top panels refer to $\frac{m_{Z^{\prime}}}{m_{N}}=2.1$ while bottom ones refer to $\frac{m_{Z^{\prime}}}{m_{N}}=5$. We fix the coupling of the $Z^{\prime}$ to $g=10^{-4}$. We also show the limit from MicroBooNE for HNL production driven by the $|U_{\mu}|$ mixing angle alone MicroBooNE:2019izn ; PhysRevD.104.055015 . For DarkQuest, we show the approved exposure benchmark of $10^{18}$ POT as well as the proposed, larger exposure of $10^{20}$ POT (green lines) in comparison with the forecasted limits using only the mixing (purple) Batell:2020vqn .
We first show results for the SBN program (SBND, MicroBooNE, ICARUS-BNB) experiments, and for the DarkQuest experiment, in their sensitivity to HNL decays with dominant mixings to either muon ($U_{\mu}$) or tau ($U_{\tau}$) flavors for two mass ratio benchmarks $m_{Z^{\prime}}/m_{N}=2.1$ and $m_{Z^{\prime}}/m_{N}=5$ in Fig. 7, where the event rate is driven by all the visible modes of the decaying HNL, including both the leptonic and hadronic final states. In calculating this number, we have adopted a 20% efficiency factor. This mimics the cuts one needs to apply to bring the background from neutrino interactions to a negligible level. In principle, the efficiency should depend on the decay channel and the mass of the HNL, but as shown in Coloma:2020lgy ; Ballett:2019bgd , using 20% is still conservative. In addition, a timing measurement of the HNL signal has been demonstrated to significantly reduce backgrounds SBND:2024vgn . The sensitivity curves are then drawn from the iso-event contours corresponding to 3 events, or a 95% confidence level of a zero Poisson background after the signal efficiency has been taken into account. Here we have fixed the gauge coupling of the $Z^{\prime}$ to be $g_{B-L}=10^{-4}$. For the SBN program experiments shown, which use the 8 GeV BNB proton beam, we do not find significant sensitivity from electron/positron bremsstrahlung, resonant production, or neutral meson decays; the primary sensitivity is due to $Z^{\prime}$ production from proton bremsstrahlung of the 8 GeV proton beam impinging on the beryllium target. We also indicate the region of the parameter space where light neutrino masses can be explained by the seesaw mechanism with the shaded band; the lower limit of this region is obtained by considering the standard seesaw formula
@@
where $\Delta m^{2}_{21}=m^{2}_{2}-m^{2}_{1}\simeq 7.5\times 10^{-5}$ eV2 is the mass difference measured by solar neutrino experiments and by KamLAND Super-Kamiokande:2016yck ; SNO:2011hxd ; KamLAND:2008dgz . The upper limit is obtained by
@@
where 0.23 eV is the current upper bound on the neutrino mass from Planck Planck:2018nkj .
@@
We draw the existing constraints derived from solely considering the mixing $U_{\alpha}$ between the HNLs and the active neutrino flavors in gray Bolton:2019pcu . These come from a variety of different experiments, including beam target or beam dump experiments that looked for HNLs being both produced and detected through their mixings; i.e., their production through meson decays to HNLs in the beam target and their subsequent decays to the same final states that we consider in this work. These include limits from PIENU Aguilar-Arevalo:2017vlf ; Aguilar-Arevalo:2019owf , PSI Daum:1987bg , E949 Artamonov:2014urb , T2K Abe:2019kgx , NA62 NA62:2020mcv , MicroBooNE MicroBooNE:2019izn , NuTeV Vaitaitis:1999wq , CHARM Vilain:1994vg ; Bergsma:1985is , NOMAD Astier:2001ck , and BEBC CooperSarkar:1985nh ; Barouki:2022bkt , for example. One can also contrast the sensitivity from only considering the $B-L$ production channels (our curves) with the mixing-only sensitivity in this fashion, like the sensitivity derived by MicroBooNE MicroBooNE:2019izn , shown in Fig. 7 by the teal dotted line. We wish to make clear that the existing constraints shown in gray do not incorporate sensitivity to the gauge boson-driven production of HNLs, but in principle, a dedicated study for each experiment could be performed to recast their bounds with the physics of a $Z^{\prime}$ in mind. This would place the comparison with our forecasted iso-event contours within the same physics model, but at present, Fig. 7 is meant to showcase how $Z^{\prime}$-driven sensitivity compares relative to the mixing-only sensitivities.
@@
Figure 8: Number of particles from all visible decays in DUNE as a function of $m_{N}$ and $\left|U_{\mu}\right|^{2}$ (left column) and $\left|U_{\tau}\right|^{2}$ (right column). Top panels refer to $\frac{m_{Z^{\prime}}}{m_{N}}=2.1$ while bottom ones refer to $\frac{m_{Z^{\prime}}}{m_{N}}=5$. We fix the coupling of the $Z^{\prime}$ to $g=10^{-4}$. The solid, dashed, and dotted lines correspond to the iso-event contours for 3, 10, and 100 events, respectively, separated by production channel (neutral meson decays, electron/positron bremsstrahlung and resonant production, and proton bremsstrahlung). We also show the forecasted DUNE-ND sensitivity to HNLs produced only from the $|U_{\mu}|$ mixing-angle driven channels (teal dotted line) Krasnov:2019kdc ; Ballett:2019bgd .
@@
Additionally, we wish to motivate the possibility of running SBND in a beam-dump or target-less mode, whereby the BNB target would be removed, and the magnetic horns turned off such that the proton beam impinges directly onto the steel beam dump. This configuration would be similar in spirit to the beam-dump run conducted at MiniBooNE MiniBooNEDM:2018cxm and the target-less configuration proposed for DUNE-ND in ref. Brdar:2022vum for their advantage in low neutrino backgrounds for BSM searches. The advantage of running in this mode lies in the much shorter distance between the HNL production site and the SBND detector, now foreshortened to only $\sim$50 meters. However, the beam power and exposure are expected to be limited in this configuration. As a realistic estimate, we take an exposure of $2\times 10^{20}$ POT and otherwise the same detector specifications. The resulting sensitivity for the SBND beam-dump mode run is shown by the dashed line in Fig. 7; we see that for only a third of the POT, the reach is roughly equivalent to the standard exposure in target mode.
@@
Lastly, we show the DarkQuest iso-event contour in green in Fig. 7 for benchmark exposures of $10^{18}$ POT and the proposed upgrade or extended run of $10^{20}$ POT (depending on the duty and plans for the Fermilab main injector beam). Note that this predicted event rate is again based on measuring all visible decay channels, both leptonic and hadronic, and we do not assume drastically different signal efficiencies between the two detection channels due to the detector capability. We find that the HNL production can also be enhanced from $B-L$ gauge boson production through proton bremsstrahlung compared to the electroweak production from neutrino mixing alone Batell:2020vqn . The $10^{20}$ POT benchmark, if it can be achieved, is especially attractive for its sensitivity to larger mass HNLs which is comparable to the DUNE-ND reach as we see next in Fig. 8.
@@
We then show results for DUNE-ND sensitivity to HNL decays with dominant mixings to either muon or tau flavors in Fig. 8. Again, we show the iso-event contours for the same two mass ratio benchmarks $m_{Z^{\prime}}/m_{N}=2.1$ and $m_{Z^{\prime}}/m_{N}=5$, this time broken up into event rate levels of 3, 10, and 100 events observed over the integrated exposure. We again adopt a 20% efficiency factor as we did for the SBN experiments. This time, the high-intensity beam and larger energies give us sensitivity to the $B-L$ driven production in both proton and $e^{\pm}$ bremsstrahlung. One can also consider the mixing-only sensitivity in this fashion for the DUNE-ND, which has been studied in refs. Krasnov:2019kdc ; Ballett:2019bgd , shown in Fig. 8 by the teal dotted line.
@@
We find that for smaller mass ratios between the $U(1)_{B-L}$ $Z^{\prime}$ mass and the HNL mass, the proton bremsstrahlung production channel at $g_{B-L}=10^{-4}$ is strong enough to probe the seesaw band. This sensitivity is roughly independent of the dominant mixing flavor, $|U_{e}|$ or $|U_{\mu}|$, but is more sensitive to different gauge coupling strengths. For example, if the $B-L$ gauge coupling is reduced by a factor of 1/2, the sensitive C.L. on $|U_{\alpha}|^{2}$ would increase by 4 since the event rate is roughly proportional to $g_{B-L}^{2}|U_{\alpha}|^{2}$ in the limit of long HNL lifetimes. Therefore we expect sensitivity to the seesaw band for gauge couplings bigger than roughly $5\times 10^{-5}$ before the event rates would be too suppressed. In comparison with the mixing-produced HNL scenario projected for DUNE-ND in teal, we note that the $Z^{\prime}$-driven production of HNLs gives us sensitivity to larger HNL masses due to the kinematic reach of proton bremsstrahlung in addition to a completely new reach to $U_{\tau}$ flavor mixings. The existing constraints in the $U_{\tau}$-$m_{N}$ parameter space are somewhat weaker than for $U_{\mu}$ and $U_{e}$ such that the $Z^{\prime}$-driven production has a much stronger relative reach.
@@ V Conclusions
By simulating the production of gauge bosons from proton bremsstrahlung, electron/positron bremsstrahlung and annihilation, and neutral meson decays, we have computed the resulting enhanced HNL flux from gauge boson decays. The flux is particularly enhanced from bremsstrahlung of the primary proton beams, a channel that sets gauged $B-L$ apart from leptophillic models, for example Capozzi:2021nmp . If the $B-L$ gauge coupling is in the neighborhood of the existing bounds, $10^{-4}-10^{-5}$, one could expect a significant enhancement to the expected HNL flux relative to the rates from standard electroweak production channels via the mixings $U_{\alpha}$. For a large enough coupling, these production rates are high enough to probe the parameter space of realistic seesaw models. This may be possible for experiments with high-intensity fluxes, as we have shown for DUNE, DarkQuest, and SBND.
@@ Acknowledgements
We thank Kevin J. Kelly for the useful feedback and discussions. We also thank Ekaterina Kriukova and Dmitry Gorbunov for their helpful comments. AT acknowledges support in part by the US Department of Energy (DOE) grant #DE-SC0010143. IMS is supported by the US Department of Energy under the award numbers DE-SC0020250 and DE-SC0020262. The work of F.C. was partially supported by the research grant number 2022E2J4RK “PANTHEON: Perspectives in Astroparticle and Neutrino THEory with Old and New messengers” under the program PRIN 2022 funded by the Italian Ministero dell’Università e della Ricerca (MUR) and by the European Union - Next Generation EU. The work of BD is supported in part by the U.S. Department of Energy Grant DE-SC0010813. The work of GG, WJ and JY is supported by the U.S. Department of Energy under Grant No. DE-SC0011686.

### v3

@@ IV Main results
Figure 7: Background-free sensitivity contours at 90% C.L. for SBND, MicroBooNE, ICARUS-BNB, and DarkQuest as a function of $m_{N}$ and $\left|U_{\mu}\right|^{2}$ (left column) and $\left|U_{\tau}\right|^{2}$ (right column). Top panels refer to $\frac{m_{Z^{\prime}}}{m_{N}}=2.1$ while bottom ones refer to $\frac{m_{Z^{\prime}}}{m_{N}}=5$. We fix the coupling of the $Z^{\prime}$ to $g=10^{-4}$. We also show the limit from MicroBooNE for HNL production driven by the $|U_{\mu}|$ mixing angle alone [41, 42]. For DarkQuest, we show the approved exposure benchmark of $10^{18}$ POT as well as the proposed, larger exposure of $10^{20}$ POT (green lines) in comparison with the forecasted limits using only the mixing (purple) [43].
We first show results for the SBN program (SBND, MicroBooNE, ICARUS-BNB) experiments, and for the DarkQuest experiment, in their sensitivity to HNL decays with dominant mixings to either muon ($U_{\mu}$) or tau ($U_{\tau}$) flavors for two mass ratio benchmarks $m_{Z^{\prime}}/m_{N}=2.1$ and $m_{Z^{\prime}}/m_{N}=5$ in Fig. 7, where the event rate is driven by all the visible modes of the decaying HNL, including both the leptonic and hadronic final states. In calculating this number, we have adopted a 20% efficiency factor. This mimics the cuts one needs to apply to bring the background from neutrino interactions to a negligible level. In principle, the efficiency should depend on the decay channel and the mass of the HNL, but as shown in [38, 11], using 20% is still conservative. In addition, a timing measurement of the HNL signal has been demonstrated to significantly reduce backgrounds [44]. In the present work, we take this signal efficiency as an assumption, as it would be subject to change given a realistic dedicated background reduction analysis by the experiments. The sensitivity curves are then drawn from the iso-event contours corresponding to 3 events, or a 95% confidence level of a zero Poisson background after the signal efficiency has been taken into account.
Here we have fixed the gauge coupling of the $Z^{\prime}$ to be $g_{B-L}=10^{-4}$, which is not yet excluded by laboratory bounds but well within potential reach of future $Z^{\prime}$ searches, e.g. LDMX and Belle-II [45], though constraints from Texono, Borexino, and CHARM II do exclude this coupling for masses below $m_{Z^{\prime}}\lesssim 100$ MeV [35, 37]. Therefore, we limit our considerations to heavier masses, or $m_{N}>50$ MeV to be conservative given the mass ratio benchmarks mentioned above. For the SBN program experiments shown, which use the 8 GeV BNB proton beam, we do not find significant sensitivity from electron/positron bremsstrahlung, resonant production, or neutral meson decays; the primary sensitivity is due to $Z^{\prime}$ production from proton bremsstrahlung of the 8 GeV proton beam impinging on the beryllium target. We also indicate the region of the parameter space where light neutrino masses can be explained by the seesaw mechanism with the shaded band; the lower limit of this region is obtained by considering the standard seesaw formula
@@
where $\Delta m^{2}_{21}=m^{2}_{2}-m^{2}_{1}\simeq 7.5\times 10^{-5}$ eV2 is the mass difference measured by solar neutrino experiments and by KamLAND [46, 47, 48]. The upper limit is obtained by
@@
where 0.23 eV is the current upper bound on the neutrino mass from Planck [49].
@@
We draw the existing constraints derived from solely considering the mixing $U_{\alpha}$ between the HNLs and the active neutrino flavors in gray [50]. These come from a variety of different experiments, including beam target or beam dump experiments that looked for HNLs being both produced and detected through their mixings; i.e., their production through meson decays to HNLs in the beam target and their subsequent decays to the same final states that we consider in this work. These include limits from PIENU [51, 52], PSI [53], E949 [54], T2K [55], NA62 [56], MicroBooNE [41], NuTeV [57], CHARM [58, 59], NOMAD [60], and BEBC [61, 62], for example. One can also contrast the sensitivity from only considering the $B-L$ production channels (our curves) with the mixing-only sensitivity in this fashion, like the sensitivity derived by MicroBooNE [41], shown in Fig. 7 by the teal dotted line.
We wish to make clear that the existing constraints shown in gray do not incorporate sensitivity to the gauge boson-driven production of HNLs, but in principle, a dedicated study for each experiment could be performed to recast their bounds with the physics of a $Z^{\prime}$ in mind. This would place the comparison with our forecasted iso-event contours within the same physics model, but at present, Fig. 7 is meant to showcase how $Z^{\prime}$-driven sensitivity compares relative to the mixing-only sensitivities. However, we argue that for many of the existing experiments the limits would not be enhanced significantly due to the presence of the $Z^{\prime}$. PSI, PIENU, and E949, for instance, all rely on the observation of $\pi$ or $K$ meson decays directly, and are more sensitive to the HNL mixing as opposed to e.g. $\pi^{\pm}\to Z^{\prime}\nu\ell^{\pm}$ decays, which are suppressed with couplings at the level of $10^{-4}$ through (if even kinematically allowed). Experiments like BEBC, PS191, T2K, CHARM, and NA62, however, constrain the heavier HNL masses and do involve powerful proton beam dumps that could potentially be sensitive to $Z^{\prime}\to NN$ production via $Z^{\prime}$ proton bremsstrahlung as we have considered. Out of these, since the T2K near detector (T2K-ND280) and CHARM are off-axis, while NA62 and PS191 make use of lower proton beam energies than BEBC, we estimate that BEBC would be the most promising of the existing experiments to gain $Z^{\prime}$-driven modifications to their limits. We performed a parameter scan using the BEBC detector and beam configurations [62, 61] and found the limit due to $Z^{\prime}$-produced HNLs in the $|U_{\alpha}|^{2}-m_{HNL}$ plane is primarily relevant for the $U_{\tau}$ mixing at lower $m_{Z^{\prime}}/m_{N}$ mass ratios; see Fig. 7, upper right. The limit would modestly exclude parameter space beyond the mixing-only limits, and in line with the aforementioned argument, the other fixed target experiments mentioned are unlikely to see limits changed as significantly by the $Z^{\prime}$-driven HNL production. For the other benchmarks we found no additional sensitivity beyond the mixing-only limits from BEBC.
@@
Figure 8: Number of particles from all visible decays in DUNE as a function of $m_{N}$ and $\left|U_{\mu}\right|^{2}$ (left column) and $\left|U_{\tau}\right|^{2}$ (right column). Top panels refer to $\frac{m_{Z^{\prime}}}{m_{N}}=2.1$ while bottom ones refer to $\frac{m_{Z^{\prime}}}{m_{N}}=5$. We fix the coupling of the $Z^{\prime}$ to $g=10^{-4}$. The solid, dashed, and dotted lines correspond to the iso-event contours for 3, 10, and 100 events, respectively, separated by production channel (neutral meson decays, electron/positron bremsstrahlung and resonant production, and proton bremsstrahlung). We also show the forecasted DUNE-ND sensitivity to HNLs produced only from the $|U_{\mu}|$ mixing-angle driven channels (teal dotted line) [63, 11].
@@
Additionally, we wish to motivate the possibility of running SBND in a beam-dump or target-less mode, whereby the BNB target would be removed, and the magnetic horns turned off such that the proton beam impinges directly onto the steel beam dump. This configuration would be similar in spirit to the beam-dump run conducted at MiniBooNE [64] and the target-less configuration proposed for DUNE-ND in ref. [65] for their advantage in low neutrino backgrounds for BSM searches. The advantage of running in this mode lies in the much shorter distance between the HNL production site and the SBND detector, now foreshortened to only $\sim$50 meters. However, the beam power and exposure are expected to be limited in this configuration. As a realistic estimate, we take an exposure of $2\times 10^{20}$ POT and otherwise the same detector specifications. The resulting sensitivity for the SBND beam-dump mode run is shown by the dashed line in Fig. 7; we see that for only a third of the POT, the reach is roughly equivalent to the standard exposure in target mode.
@@
Lastly, we show the DarkQuest iso-event contour in green in Fig. 7 for benchmark exposures of $10^{18}$ POT and the proposed upgrade or extended run of $10^{20}$ POT (depending on the duty and plans for the Fermilab main injector beam). Note that this predicted event rate is again based on measuring all visible decay channels, both leptonic and hadronic, and we do not assume drastically different signal efficiencies between the two detection channels due to the detector capability. We find that the HNL production can also be enhanced from $B-L$ gauge boson production through proton bremsstrahlung compared to the electroweak production from neutrino mixing alone [43]. The $10^{20}$ POT benchmark, if it can be achieved, is especially attractive for its sensitivity to larger mass HNLs which is comparable to the DUNE-ND reach as we see next in Fig. 8.
@@
We then show results for DUNE-ND sensitivity to HNL decays with dominant mixings to either muon or tau flavors in Fig. 8. Again, we show the iso-event contours for the same two mass ratio benchmarks $m_{Z^{\prime}}/m_{N}=2.1$ and $m_{Z^{\prime}}/m_{N}=5$, this time broken up into event rate levels of 3, 10, and 100 events observed over the integrated exposure. We again adopt a 20% efficiency factor as we did for the SBN experiments. This time, the high-intensity beam and larger energies give us sensitivity to the $B-L$ driven production in both proton and $e^{\pm}$ bremsstrahlung. One can also consider the mixing-only sensitivity in this fashion for the DUNE-ND, which has been studied in refs. [63, 11], shown in Fig. 8 by the teal dotted line.
@@
We find that for smaller mass ratios between the $U(1)_{B-L}$ $Z^{\prime}$ mass and the HNL mass, the proton bremsstrahlung production channel at $g_{B-L}=10^{-4}$ is strong enough to probe the seesaw band. This sensitivity is roughly independent of the dominant mixing flavor, $|U_{\mu}|$ or $|U_{\tau}|$, but is more sensitive to different gauge coupling strengths. For example, if the $B-L$ gauge coupling is reduced by a factor of 1/2, the sensitive C.L. on $|U_{\alpha}|^{2}$ would increase by 4 since the event rate is roughly proportional to $g_{B-L}^{2}|U_{\alpha}|^{2}$ in the limit of long HNL lifetimes. Therefore we expect sensitivity to the seesaw band for gauge couplings bigger than roughly $5\times 10^{-5}$ before the event rates would be too suppressed. In comparison with the mixing-produced HNL scenario projected for DUNE-ND in teal, we note that the $Z^{\prime}$-driven production of HNLs gives us sensitivity to larger HNL masses due to the kinematic reach of proton bremsstrahlung in addition to a completely new reach to $U_{\tau}$ flavor mixings. The existing constraints in the $U_{\tau}$-$m_{N}$ parameter space are somewhat weaker than for $U_{\mu}$ and $U_{e}$ such that the $Z^{\prime}$-driven production has a much stronger relative reach.
@@ V Conclusions
[…]
@@ Acknowledgements
[…]